March BG Com Night 5, March 10-11 2009 MST

Observers: J. Hill, A. Rakich (Tucson) plus J. Storm (Potsdam)
Lucifer support: N. Ageorges (Garching)
Telescope Operator: A. Ceranski
SW Support: N. Cushing, C. Biddick (Tucson)
Telescope Support: J. Little

Summary:

High humidity and distant thunderstorms keep us closed for the first two hours.

Then we got into some bizarre state where the pointing and tracking are strange.

The mirror support air vent valve provided a plumbing diversion near midnight.

After reopening, we found with LUCIFER that the telescope pointing was off by 9.5 arcmin in Elevation. We still don't know why.

We verified that the WFS hotspot on the guide camera is just where it was supposed to be, and we measured the new position of the pupil center on the WFS CCD.

Then we got a good start on taking transformation data for Jesper.

Approaching clouds and radar echoes (snow) forced us to close at 4AM.

Details (Times in MST)

Closed for high humidity early.

Open

20:00 Open Vent Doors

20:05 Open Dome, T=-3 degC, D=-5.5 degC

20:13 restart GCSL after removing starOverlay and restoring sextactor centroiding.

20:15 pointto M5_0258 gs=0 ACQUIRE PAM-

20:20 Moved IE and CA to -13, -20 from Saturday, but still don't see any star.

20:22 cl < home$spiral.cl found star at 70,70. Added 0.6 mm Zglobal to focus.

But stars are drifting......... ???? at a rate of 20 arcsec/minute. We were at AZ=218 EL=55.

pointto found a star (random?) at offset 30 30 This led to IE=-35 CA=-45, but it is a moving target.

Called Michele and Dave, both PCS and MCS are the same as Sunday morning

21:15 pointto M5_0444 ACQUIRE PAM-

21:20 pointto M5_0444 ACQUIRE

21:21 cl < home$spiral.cl

GPS receiver is locked.

21:38 slew to M42 - PCSGUI shows M42 in the wrong place. PCSGUI display doesn't work correctly for negative declinations, but that is only an incidental problem along the way.

AZ and EL agree with xephem.

Preset to M44

lbtmu102 and jet both have the correct date and time.

22:40 Killed and restarted MCS and MCSPU.

22:45 Preset to M42 05:35:20 -5:23:12

cl < home$spiral.cl without success

23:12 pointto M5_0444 TRACK

Tom arrives from the airport. Michele is on the phone. Chris manages to revive the MAT.

Both this star and M42 at least show up on the wide-field MAT. MAT seems to indicate that we aren't more than 10 arcmin off.

/lbt/mat/bin/MATGUI is the secret MAT incantation. Start a ds9 window beforehand.

cl < home$spiral.cl This spiral covers a 4 arcminute square box, and we don't see the star.

23:34 Switch back to the B22 default PCS. Stop and restart.

23:36 pointto M5_0444 TRACK, pointing is the same.

23:40 Back to the newest PCS. This is the PCS that was working OK last Sunday morning.

23:43 preset Polaris 02 31 49.084 +89 15 50.79

We saw Polaris on the MAT (at a position consistent with where we saw M42). Then it disappeared. Clouds in the north?

23:50 Primary mirror panics. Lost air pressure. But why did we lose the star on the MAT when the primary panic'd? One of the M1 dewpoint valves is chattering.

00:01 Close for rising humidity while we are sorting out this air problem.

00:20 John L. and Aaron have the mirror valve under control. So now we are raising the mirror. They closed the cross-over valve, so this chattering valve won't bother us anymore tonight.

00:30 Resetting the mount encoders again.

00:35 Reopening enclosure with humidity, 86.5%

00:41 John crashes the Polycom box, that is a sign of a bad night.

Tried looking for Polaris, but there seems to be a cloud there.

01:07 pointto M5_0444 TRACK

Found in LUCIFER at IE=565 CA=0. This is 364,371 pixels on WF-MAT.

pointto M5_04444 gs=0 ACQUIRE

01:34 ACTIVE - UMAC error

01:35 ACTIVE - succeeds but GCS does recognize the comatic star in the center of the acquisition image, acquire00004.

01:37 pointto BS9152 gs=0 ACTIVE - star verification failed.

01:42 Edit lbt.conf to use internal centroids again. restart GCSL.

Verify Hotspot and WFS Pupil on sky (estimated 2 hours)

01:43 pointto BS9152 gs=0 ACTIVE - succeeds! We have a 1 arcsec image with loads of coma. We see spots on the WFS image, so the star is close to the correct hotspot. In only a few iterations we are already down to 0.6 arcsec FWHM.

01:50 Block guiding corrections on PCS, with GCS still taking guide images, while Andrew takes WFS images with IDL.

02:00 Confirmed that we are well-centered on the hotspot in the IE direction (presently upper left to lower right on the guide image)

02:10 Confirmed that we are well-centered on the hotspot in the CA direction. Last image was wfsc00021. in 090311w/ Conclude that the old hotspot position in GCS is good.

02:11 Now guiding again and collimating with IDL IDL auto centering is giving us 401.7, 220.9 for the pupil center. Editted AIP_L.cfg and the corresponding IDL file to record this center.

We've emailed Doug to see if we need to shift the lenslet edge parameters by one lenslet in X. (We did, and we fixed it the following day.)

02:44 Editted PCSInstrument.conf to use LEFTZEROPOINT = -187.703 which is DJT's carefully analyzed value from the 20090214 data.

02:48 Editted lbt.conf to change focal length to DJT's improved value of 123421.4 mm. Why does lbt.conf have two focal lengths?

Transformation data - step 1

Nancy is on the telephone. Jesper and Aaron are on Polycom. Jesper is collecting data with /home/agwuser/getdat which has one argument which is the star number to be logged.

03:01 pointto BS9177 gs=0 ACTIVE - starting at EL=36 deg with 0.8 arcsec FWHM.

03:06 pointto BS9177 gs=0 TRACK PA=0 - luci_20090311_0002

03:11 pointto BS9177 gs=0 TRACK PA=180 - luci_20090311_0003

03:13 pointto BS9177 gs=0 TRACK PA=90 - luci_20090311_0004

03:16 pointto BS9177 gs=0 TRACK PA=270 - luci_20090311_0005

Transformation data - step 2

03:19 pointto BS9177 gs=0 ACQUIRE IE=+562 CA=+11

03:23 pointto BS9177 gs=0 ACTIVE PA=0 (no image as probe blocks field) 0.6 arcsec FWHM

03:28 pointto BS9177 gs=43 ACTIVE PA=0 - luci_20090311_0006

03:31 pointto BS9177 gs=49 ACTIVE PA=0 - (no star in acquire00005) But note that GCS is not updating the displayed filename when it displays the acquisition image.

03:32 pointto BS9177 gs=53 ACTIVE PA=0 - (star was to faint for internal centroiding)

03:34 pointto BS9177 gs=57 ACTIVE PA=0 - out of range of probe

03:36 pointto BS9177 gs=19 ACTIVE PA=0 - luci_20090311_0007

Star is quite jumpy at the start of guiding. What is causing this? The first guide image after the acquisition offset is exactly on target, but the second guide image jumps off by 1 arcsec, and then the third one is OK. After that guiding is normal. ???? This happened several times between 03:30 and 03:55 MST.

03:39 pointto BS9177 gs=46 ACTIVE PA=0 - luci_20090311_0008

03:42 pointto BS9177 gs=47 ACTIVE PA=0 - luci_20090311_0009

03:45 pointto BS9177 gs=58 ACTIVE PA=0 - luci_20090311_0010

03:48 pointto BS9177 gs=52 ACTIVE PA=0 - luci_20090311_0011 (too dim for WFS)

03:51 pointto BS9177 gs=45 ACTIVE PA=0 - wait for OSS in-position timed out

03:53 pointto BS9177 gs=45 ACTIVE PA=0 - luci_20090311_0012

03:56 pointto BS9177 gs=50 ACTIVE PA=0 - luci_20090311_0013

03:59 pointto BS9177 gs=31 ACTIVE PA=0 - luci_20090311_0014

Close

04:02 Close for approaching radar echoes and puffy clouds.

04:30 It is starting to snow on Mt. Graham, so we are done for tonight.

-- JoarBrynnel - 02 Mar 2009
Topic revision: r9 - 27 May 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback